college algebra
solve each system by the addition method. Determine whether the equation are independent, dependent, or inconsistent.
x-y=3
-6x+6y=17

Answer by richwmiller(17219)

multiply the first by 6
6x-6y=18
-6x+6y=17
add
0=35
They are parallel. They never cross. There are no solutions. So they are inconsistent.